### v1

# Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures

Abstract
Semiconductor–superconductor hybrid devices have been proposed as promising platforms for detecting and analyzing Majorana zero modes, which find applications in topological quantum computing. In this work, we solve the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe the normal and superconducting properties of a PbTe/Pb heterostructure. We resolve a proximity–induced superconducting gap on the PbTe side. The hybridization between PbTe and Pb causes the emergence of a soft Bardeen–Cooper–Schrieffer–like superconducting gap. We compute the anomalous charge density in real space, estimating its decay length and showing that the pairing potential is anisotropic, which is a necessary condition for unconventional superconductivity. Contrary to the models that predict Majorana zero modes in these interfaces, we find a significantly large Schottky barrier in the normal state preventing the emergence of zero modes. Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their viability for Majorana zero modes applications.
keywords:
superconductivity, semiconductor-superconductor heterostructure, BdG equations
1 Introduction
PbTe/Pb nanowires are a promising material platform to host Majorana Zero Modes (MZMs) due to their resilience to charged impurities, high electron mobility, and excellent crystalline quality  1, 2. The intrinsic properties of PbTe (narrow bandgap and large Landé g-factor) combined with the superconducting gap ($\Delta\sim$ 3meV) and relatively high critical temperature ($T_{C}\sim$ 7K) of Pb, make these materials advantageous compared to conventional InAs-Al and InSb-Al nanowires 3, 4.
To host MZMs, a semiconductor–superconductor (SM–SC) interface must meet several key requirements 5, 6: (i) strong spin-orbit coupling and significant Zeeman energy gap ($V_{Z}$) in the semiconductor, (ii) $s$–wave superconductor, (iii) proximity induced superconductivity, and (iv) the heterostructure chemical potential must lie within the SM’s Zeeman gap and the SC gap. When the applied Zeeman field exceeds the critical value $V_{C}=\sqrt{\Delta^{2}+\mu^{2}}$ the system transitions into a topological phase, hosting MZMs at its ends 7, 8, 9.
Cole et al. 10 developed a model to describe the density of states (DOS) in a generic nanowire and investigated how the SC–coupling strength and the height of the interface barrier affect the shape of the SC gap. They identified a hard U–shaped superconducting gap as indicative of a strong SC–coupling regime, where the superconductor overshadows the semiconductor’s electronic properties. Conversely, a soft Bardeen–Cooper–Schrieffer–like gap corresponds to a weak SC–coupling regime, where low-energy states of the HS are predominantly of semiconductor character. Weak SC–coupling is, therefore, desirable for achieving topological superconductivity 11, 12.
In this work, we investigate from first–prinicples proximity effects in a PbTe/Pb heterostructure (HS), focusing on the conditions required for MZMs to emerge. All our findings are obtained using the recently implemented SIESTA-BdG method 13, 14, which solves the Kohn-Sham Density Functional Theory and Bogoliubov–de Gennes equations to describe on the same footing the normal and superconducting properties. We discuss how the electrostatic environment at the interface is critical for the emergence of MZMs. We find significant band bending at the interface, leading to charge transfer from PbTe to Pb, and affecting the position of the chemical potential. In the superconducting state we observe a soft BCS–like SC gap, indicating a weak SC–coupling between PbTe and Pb. We predict a proximity–induced superconducting gap in PbTe and a softening of Pb SC gap (with respect to bulk). We show that the superconducting anomalous charge density ($\chi(\mathbf{r})$) emerging at the interface is anisotropic. Our findings are robust against lattice deformations (strain) and applied electric field. In particular, the induced SC gap at the interface is resilient under lattice strain.
Figure 1: (a) Side view of the relaxed (4, -5) PbTe/Pb heterostructure stacked along the [001] direction. (b) Average electrostatic potential ($\Delta V_{ave}$) across the interface with a schematic of the energy alignment between a semiconductor and a metal. (c) Schematic energy alignment between a semiconductor and a superconductor, representative of PbTe/Pb. The optimal alignment for the emergence of Majorana Zero Modes occurs when the semiconductor gap and the superconductor gap align in energy.
2 PbTe/Pb Heterostructure: Normal state
[…]
In the normal state, the bands (Figure 2a) cross the Fermi level, resulting in a metallic heterostructure with visible band splitting (mostly near the $M$ point) due to hybridization between PbTe and Pb states. To prove this, we remove the interaction between PbTe and Pb by setting their coupling to zero in the Kohn Sham Hamiltonian. The resulting band structure (Figure 2b) does not exhibit any splitting. We highlight the PbTe contribution to the electronic structure with dark red points in (Figure 2b), proving that the states near the Fermi energy predominantly belong to the semiconductor.
Figure 2: (a) PbTe/Pb normal state band structure. The bands cross the Fermi level, resulting in a metallic heterostructure. (b) Normal state bands with the coupling between the PbTe and Pb sides of the heterostructure set to zero. The color intensity represents the contribution from the PbTe orbitals.
The interaction between the two materials leads to variation in chemical bonds and relative atomic positions near the interface, along with charge transfer from PbTe to Pb. Using Mulliken charge analysis 15, we compute the variation of charge $\Delta Q$ on PbTe ($Q_{PbTe}$) and Pb ($Q_{Pb}$) in the HS with respect to their freestanding nominal values. We find a depletion of charge on PbTe indicating an electron transfer from the PbTe to the Pb side of the HS (Supporting Information, FigureS11).
When a metal comes in contact with a semiconductor, their energy level align, creating a potential energy barrier (Schottky barrier) at the interface due to the difference in their work functions (Figure 1b). We calculate from first-principles the average electrostatic potential ($\Delta V_{ave}(z)$)16 , namely the spatial average of all electrostatic interactions in the system. The average is performed layer-by-layer, and it quantifies the band alignment in heterostructure. It is desirable to have a $\Delta V_{ave}\sim 0$ eV so that the electronic states participating in the superconducting proximity effect are close to the SOC splitted bands of PbTe at $L$ (Figure 1d) and Supporting Information, Figure S8]. In this case, the requirements for emergence of MZMs would be satisfied 7. We predict, instead, a significant difference in the average electrostatic potential across the interface $\Delta V_{ave}\sim 1.2$ eV, with the PbTe side being higher in energy than the Pb side (Figure 1b).
We studied the robustness of these findings against strain and the presence of an external electric field. Our results indicate that $\Delta V_{ave}$ and the normal state electronic DOS near the Fermi level are not significantly affected by varying strain (Table 1 and Figure 3a). We compute the excess charge $\Delta Q$ and observe that it slightly increases by increasing strain on PbTe (Table 3), substantiating the conclusion that the electrostatic environment close to the interface does not change under lattice deformations. Furthermore, the shape and magnitude of the DOS is not affected by the electric field (Figure 3b). The magnitude of the applied electric fields are all considerably above the electrical breakdown point of bulk PbTe which can be estimated to be $6.4\times 10^{-6}$ V/Å 17, while its sign and direction are chosen to offset the difference in electrostatic potential at the interface $\Delta V_{ave}$.
Figure 3: PbTe/Pb normal state DOS for different (a) strain and (b) applied electric field across the interface, represented with different colors.
Table 1: Difference in average electrostatic potential $\Delta V_{ave}$ across the interface of the PbTe/Pb HS for different strain. The (4, -5) HS is highlighted in bold.
strain (PbTe/Pb) [%] $\Delta V_{ave}$ [eV]
1.0%/-7.7% 1.36
2.0%/-6.8% 0.84
3.0%/-5.9% 0.93
4.0%/-5.0% 1.20
5.0%/-4.0% 1.08
6.0%/-3.2% 1.10
7.0%/-2.3% 1.17
8.0%/-1.3% 1.22
9.0%/-0.4% 1.30
3 PbTe/Pb heterostructure: Superconducting state
Figure 4: Superconducting DOS (SC–DOS, $\rho(\varepsilon)$, black continuous line) and anomalous DOS (ADOS, $\chi(\varepsilon)$, dashed green line) for the (4, -5) PbTe/Pb HS. Two distinct energy ranges are resolved within the soft BCS–like SC gap: the difference between the two main coherence peaks $\Delta$ and the induced SC–gap $\Delta_{ind}$, where the SC–DOS vanishes. The dashed black line serves as a reference for zero DOS. (a) SC–DOS projected on individual PbTe (b) and Pb (c) layers, and for the PbTe/Pb HS for diffrent strain values (d).
We used the SIESTA-BdG method to compute the superconducting DOS of bulk Pb (Supporting Information) and PbTe/Pb heterostructures. We demonstrate that a proper initialization of the superconducting pairing potential is essential to avoid unphysical results (Supporting Information Figure S10). Moreover, we demonstrate how the SIESTA-BdG method can be used to investigate various pairing mechanisms and their effects on the superconducting density of states (SC–DOS). Computational details and geometrical aspects of the HS are presented in the Supporting Information Section S6.
Bulk Pb exhibits a hard U–shaped gap of $\Delta\sim 3.0$ meV. The HS, instead, exhibits a soft U–shaped gap with coherence peaks $\Delta=2.0$ meV far apart and an induced gap $\Delta_{ind}=0.6$ meV (Figure 4a). We find that the anomalous density of states (ADOS, $\chi(\varepsilon)$) is dominated by the singlet components (Cooper pair with total spin zero), while the triplet components (Cooper pair with total spin one) are negligible. The shape of the total DOS is determined by the induced gap in the PbTe side of the heterostucture, as indicated by the projected density of states (Figure 4b). The low energy states inside the SC gap are predominantly of PbTe character, hence the HS is in the weak coupling regime 10. The hybridization between Pb and PbTe “poisons” the SC gap, leading to a soft BCS–like superconducting density of states (SC–DOS $\rho(\varepsilon)$). The Pb side of the heterostructure retains its U-shape but the coherence peaks are less sharp and smaller in magnitude than bulk Pb[Figure 4.(c)]. In our simulations, the soft–BCS gap at the SM–SC interface is due to structural relaxation of the atomic structure and states hybridization, as we do not include magnetic, thermal, or dissipative broadening effects 18. We find that strain only slightly affects the shape of the SC DOS in the HS (Figure 4d).
Figure 5: (a) Anomalous density of states $\chi(\mathbf{r})$ along the z-axis in real space. Vertical dashed lines represent the layer position in the PbTe (olive) and Pb (black) regions of the heterostructure. (b) 3D visualization of the anomalous DOS near the interface region. The isosurface level, indicate regions where $\chi(\mathbf{r})$ is constant. To enhance visualization, the isosurface on the PbTe side of the heterostructure (blue) is set to a value 10 times smaller than the Pb side (red).
From the computed anomalous charge density $\chi(\mathbf{r})$, we explain the SC–gap dependence from the position along the heterostructure. On the Pb side, $\chi(\mathbf{r})$ oscillates around a constant value with maxima at the Pb atomic positions (Figure 5a). The oscillations are likely due to the finite size of the simulated system. On the PbTe side, $\chi(\mathbf{r})$ decays from the interface towards the surface with a superconducting coherence length $\xi\sim 14$. The coherence length ($\xi$) was computed by fitting an exponential function to $\chi(\mathbf{r})$.
We visualize the anomalous charge density in (Figure 5b), where the isosurface level indicate regions with constant $\chi(\mathbf{r})$. To enhance visualization, the isosurface on the PbTe side of the heterostructure [blue regions in Figure 5.(b)] is set to a value 10 times smaller. The anomalous charge density penetrates in the PbTe region, it is enhanced at the interface, and it is highly anisotropic. On the Pb side of the HS, the anomalous density is homogenous in the central region (layers 21–22) similar to bulk Pb, while it deforms around the interface and the vacuum region. The anomalous charge density is directly proportional to the superconducting pairing potential 13. Therefore, the pairing potential in the HS is not isotropic. The anisotropy of the pairing potential constitute a signature of unconventional superconductivity.
4 Conclusion
We used a combined Kohn–Sham and Bogoliubov–de Gennes approach to investigate the semiconductor–superconductor (SM–SC) heterostructure composed of PbTe and Pb, focusing on the interfacial properties and the potential applicability of this system for detecting Majorana Zero Modes (MZMs) 19, 9, 7. We discussed the (normal state) metallization of the heterostructure driven by strong hybridization between PbTe and Pb states at the interface. We predict a significant difference in electrostatic potential at the interface, which poses a challenge in achieving the necessary energy alignment between the SM Zeeman gap and SC gap. These findings remain robust against strain and electric field.
In the superconducting state, a soft BCS–like superconducting gap is computed, indicating a weak superconducting coupling between the two systems 10. Moreover, we predicted proximity–induced superconductivity in the PbTe region of the heterostructure. We resolved the superconducting density of states layer by layer, revealing that the induced SC gap is enhanced near the interface, with a softened SC gap on the Pb side. We analyzed the shape and decay of the anomalous charge density $\chi(\mathbf{r})$, revealing an anisotropic pairing potential. We observed a weak superconducting coupling regime at the interface, which prevents the superconductor to overscreen the HS properties. The weak regime favours the device’s tunability with gate and bias voltages, creating a route to get the desirable energy alignment between the two materials. Our findings strengthen the understanding of the physics governing PbTe/Pb hybrid devices and their potential application in hosting Majorana Zero Modes (MZMs).
Acknowledgments
The authors acknowledge the fruitful discussion with Erik Bakkers, Arnold H. Kole and Nils Wittemeier. ZZ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). RR and ZZ acknowledge financial support from Sector Plan Program 2019-2023. This work was sponsored by NWO-Domain Science for the use of supercomputer facilities. We also acknowledge that the results of this research have been achieved using the Tier-0 PRACE Research Infrastructure resource Discoverer based in Sofia, Bulgaria (OptoSpin project id. 2020225411). This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No 101130384 (QUONDENSATE).
Supporting Information for:
Ab initio study of Proximity-Induced Superconductivity in PbTe/Pb heterostructures
R. Reho, Andrés R. Botello-Méndez, and Zeila Zanolli
The PDF file includes:
Normal and superconducting properties of bulk Pb and PbTe, theoretical methods, details concerning the initialization of the superconducting pairing potential, geometrical aspects of the heterostructures, and supplementary figures.
S5 Supplementary note 1: Computational methods
To model the PbTe/Pb interface, we computed its electronic structure using density functional theory (DFT) as implemented in the SIESTA code 20, 14, including spin-orbit coupling and the PBE exchange-correlation functional 21. We employed optimized norm-conserving Vanderbilt pseudopotentials (ONCVPSP 22) in PSML format 23 from PseudoDojo database 24. The Kohn-Sham equations were solved using standard double-zeta polarized basis sets of localized atomic orbitals on a $60\times 60\times 1$ Monkhorst-Pack grid. Real space integral were evaluated on grids with a cut-off energy of 2000 Ry for the SC state calculations and 500  Ry for the normal state ones. The Fermi-Dirac occupation function was smoothed using an electronic temperature of 0.1 meV, to ensure that the broadening of the electronic states is well below the size of the superconducting gap. All crystal structures were relaxed using the FIRE algorithm with a force threshold of 0.005 eV/Å and a maximum stress tolerance of 0.5 meV/Å3.
The superconducting properties of PbTe/Pb were modeled using the SIESTA–BdG method 13. In this approach a semi-empirical superconducting pairing potential is introduced on top of the self-consistent Kohn-Sham Hamiltonian describing the normal state. We used the fixed–$\Delta$ method with the superconducting pairing potential initialized in real space as a touching spherical hardwells with radius $\mathbf{r}=1.58$ and strength $\bar{\Delta}=1.50$ meV around Pb atoms. We disregard the singlet–triplet mixing effect 25 as the computed contributions from triplet states are considerably lower in magnitude than the singlet ones.
Figure S6: Orbital (a) and spin (b) projected bands structure of bulk Pb in the normal state. In (a) we observe that the electronic states crossing the Fermi level are $p$ and $d$ orbitals. In (b) we observe a spin flip at $W$ for $S_{x}$ and $S_{y}$ spin orientations.
S6 Supplementary note 2: Geometrical aspect of the HS
We constructed the PbTe/Pb HS, stacking 15 layers slab of PbTe cell and 11 layers of Pb along the [001] direction. The HS extends infinitely in the x and y directions. Periodic replicae are separated by 26 Å  vacuum along the z–axis. The PbTe slab was constructed by rotating the in-plane conventional cell by 45 degrees and reconstructing its primitive cell. The Pb slab was constructed by repeating the conventional cell along the [001] direction 12 times and removing one layer to achieve 11 layers of Pb. Farey sequences 26 were employed to identify commensurate cells for the full heterostructure. This method inherently introduces strain in the system. We considered several strained configurations of the PbTe/Pb heterostructure. The strain on each side of the heterostructure is defined as:
$\text{strain}=\frac{\mathrm{a}_{\mathrm{hs}}-\mathrm{a}_{\mathrm{rlx,i}}}{\mathrm{a}_{\mathrm{rlx,i}}}\quad\mathrm{i=Pb,PbTe}$ (S1)
where $\mathrm{a}_{\mathrm{hs}}$ is the lattice parameter of the heterostructure while $\mathrm{a}_{\mathrm{rlx,i}}$ is the relaxed lattice parameter of the constituent materials.
S7 Supplementary note 3: Bulk Pb and bulk PbTe
We present the crystal structure and electronic properties of bulk Pb and PbTe, emphasizing the orbitals’ contribution to the electronic states near the Fermi level. The interaction and hybridization between these states are critical for the proximity–induced superconductivity effect at the PbTe/Pb interface.
Figure S7: (a) SC–DOS (continous lines) and ADOS (dashed lines) of bulk Pb computed with the solution methods: non SCF-BdG (black), Fixed-$\Delta$ (blue), and Full SCF-BdG (red). The Full SCF-BdG method reveals a small splitting of the main coherence peaks. (b) Orbital contribution to the DOS using the Full SCF-BdG method, showing that the main contributions from $p$ and $d$ orbitals, as indicated by the orbitals crossing the Fermi level. Similar results are found using the other two solution methods.
S7.1 Bulk Pb
[…]
We modeled the superconducting state using the SIESTA-BdG code 13, employing three solution methods: non SCF-BdG, Fixed-$\Delta$, and full SCF-BdG. All three solution methods predict a conventional U-shaped SC gap of $\sim$3 meV (Figure S7). We initialize the pairing potential in the superconducting strength representation (on a real-space grid) as touching spherical hardwells with radius $\mathbf{r}=1.58$. In the non SCF–BdG and fixed–$\Delta$ methods, the initial strength of the superconducting pairing potential is $\bar{\Delta}=1.50$ meV. In the full SCF-BdG method, instead, the initial value of the superconducting coupling is $\lambda(\mathbf{r})=80.3$ eV. We computed the orbital projected superconducting density of states (Figure S7) and conclude that Pb $p$ and $d$ are the dominant electronic states close to the Fermi level, and hence relevant for the superconducting properties.
S7.2 Bulk PbTe
[…]
Table 2: Electronic gap ($\Delta E_{L}$) of PbTe as a function of its in-plane lattice parameter. The configuration that matches the value of the lattice parameter for the (4,-5) PbTe/Pb HS is highlighted in bold.
Lattice parameter [Å] $\Delta E_{L}$ [eV]
4.39 0.251
4.44 0.224
4.49 0.147
4.53 0.069
4.58 0.023
4.63 0.058
4.67 0.107
4.72 0.149
4.76 0.181
Figure S8: Electronic band structures for PbTe bulk. (a) Orbital projected band structure. The inset shows an energy region close to the Fermi level. (b) Spin projected band structure. Positive and negative spins are depicted in red and blue, respectively.
Figure S9: Projected DOS of bulk PbTe. Near the Fermi level, the valence band is predominantly influenced by Te $p$ orbitals, whereas the conduction band shows a substantial contribution from Pb $p$ orbitals
S8 Supplementary note 4: Initialization of the Pairing Potential and Its Impact on the Superconducting Properties
In the SIESTA-BdG method one can employ different methods for solving the BdG equations and different approaches for initializing the pairing potential $\Delta$ 13. We tested multiple initializations of the pairing potential. In the main text, we report results obtained within the superconducting strength representation for all simulations.
We compared the superconducting strength and orbital representation (intra–orbital coupling or full–overlap) for the PbTe/Pb heterostructure (Figure S10). The intra–orbital coupling lacks the interactions between different orbitals that can be captured by the superconducting strength representation, making it more localized. As a result, the intra–orbital coupling misses the hybridization and coupling between states belonging to the PbTe and Pb sides of the heterostructure, leading to a U–shaped DOS. However, when we excessively couple states belonging to different orbitals and atoms (full–overlap), the superconducting gap closes. Therefore, initializing the pairing potential in real–space as touching sphere hardwells is a physically intuitive and reasonable approach.
Figure S10: The SC–DOS, $\rho(\varepsilon))$, of the PbTe/Pb heterostructure is computed using the superconducting strength representation method (solid black line) and the orbital representation method. In the orbital representation, we evaluate intra-orbital coupling (solid orange line) and full-overlap coupling (dashed blue line). When pairing occurs between electrons and holes within the same orbital on the same atom (intra–orbital), a conventional U-shaped gap emerges. In contrast, when superconducting coupling involves interactions between different orbitals across different atoms (superconducting strength representation or full–overlap), the hard superconducting gap becomes softened.
S9 Supplementary note 5: Mulliken charge analysis
Table 3: Excess of charge $\Delta Q$ computed for different strain. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. We compare the excess of charge on the PbTe side ($\Delta Q_{PbTe}$) of the heterostructure, with the one on the Pb side ($\Delta Q_{Pb}$). The (4, -5) HS is highlighted in bold.
strain (PbTe/Pb) [%] $\Delta Q_{PbTe}$ [1/e] $\Delta Q_{Pb}$ [1/e]
1.0%/-7.7% -0.096 0.096
2.0%/-6.8% -0.098 0.098
3.0%/-5.9% 0.101 0.101
4.0%/-5.0% -0.103 0.103
5.0%/-4.0% -0.107 0.107
6.0%/-3.2% -0.106 0.106
7.0%/-2.3% -0.109 0.109
8.0%/-1.3% -0.114 0.114
9.0%/-0.4% -0.119 0.119
Figure S11: Mulliken charge analysis for PbTe/Pb. Layer-by-layer excess charge. Positive (negative) values of $\Delta Q$ denote the acceptor (donor) character of the layer. The red dashed line denote the separation between the PbTe and Pb side of the heterostructure. We notice that at the interface there is a depletion of charge on the PbTe side of the heterostructure, while the opposite is true on the Pb side. The sum of the excess charge on the PbTe and Pb side $\Delta Q$ is reported in the plot.